H. Cochran, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2,410,000. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,710,000 in annual sales, with costs of $1,730,000. Assume the tax rate is 23 percent and the required return on the project is 9 percent. What is the project’s NPV? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,410,000/3)=$803,333.333

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(2,710,000-1,730,000)(1-0.23)+(0.23*803,333.333)

=$939366.667(Approx).

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=$939366.667[1-(1.09)^-3]/0.09

=$939366.667*2.53129467

=$2377813.83

NPV=Present value of inflows-Present value of outflows

=$2377813.83-$2,410,000

=($32186.17)(Approx)(Negative).